# One Example Shown, Many Concepts Known! Counterexample-Driven Conceptual Reasoning in Mathematical LLMs

### v1

@@ 1 Introduction
As illustrated in Figure 1, for human mathematics learning, “example-based” learning is a more important strategy than drill-based learning. In particular, for mathematical proofs, “proof by counterexamples” is an indispensable approach. Inspired by the idea that counterexample-driven proofs better reflect a deep understanding of mathematical concepts, we propose CounterMATH, a counterexample-based mathematical reasoning benchmark. CounterMATH is designed to evaluate LLMs’ ability to distinguish subtle differences between mathematical terms and properties at university-level by providing examples. Specifically, we collect 1,216 statement-rationale pairs from mathematical textbooks, focusing on disproving certain statements under unusual conditions using counterexamples. In terms of difficulty, CounterMATH covers advanced mathematical knowledge similar to PutnamBench (Tsoukalas et al., ) and Putnam-AXIOM (Gulati et al., 2024), both of which assess the depth of mathematical understanding in LLMs.
@@ 4 Benchmark Settings
We are testing following large language models with CounterMATH. For open-weight models, we use Deepseek-Math-7B-RL (Shao et al., 2024), Eurus-2-7B-PRIME (Cui et al., 2025), Qwen2.5-Math-7B/72B-Instruct (Yang et al., 2024), NuminaMath-7B-TIR (Li et al., 2024b), InternLM2-Plus-7B/20B/Mixtral8x22B (Ying et al., 2024), Abel-7B/13B/70B (Chern et al., 2023), WizardMath-7B/70B (Luo et al., 2023), Mathstral-7B [^3] https://mistral.ai/news/mathstral/, MetaMath-Mistral-7B (Yu et al., 2023), Xwin-Math-7B/13B/70B (Li et al., 2024a), Rho-math-7b-interpreter (Lin et al., 2024b), MAmmoTH2-7B/8x7B-Plus (Yue et al., 2024), QwQ-32B-Preview [^4] https://qwenlm.github.io/blog/qwq-32b-preview/. For proprietary models, we use GPT-4o, OpenAI o1555https://cdn.openai.com/o1-system-card-20241205.pdf, Qwen-Max [^6] https://www.alibabacloud.com/help/en/model-studio/developer-reference/what-is-qwen-llm. The selection of baseline models tries to cover as many representative models as possible, which considers various perspectives, including data processing, training paradigms, base models, and developers (i.e., academia or industry). The detailed summary of open-weight baseline models is in Appendix B.
@@
5 Conceptual Finetuning
5.1 Training Data Engineering Framework
Filter-based Data Collection
@@ 5.2 Training and Evaluation Details
For model training, we select Qwen-2.5-Math-7B-Instruct, an open-source model known for its strong mathematical reasoning capabilities and general applicability. After training, we evaluate the model on both our CounterMATH and several out-of-distribution (OOD) benchmarks, such as MATH(Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021). We aim to demonstrate not only improved performance on the CounterMATH but also enhanced generalization to OOD benchmarks, thereby validating the effectiveness of our insight in improving the model’s overall conceptual reasoning capabilities by examples.
@@ 6.1 Evaluation Results without Finetuning
The performance on the automatic evaluation metric F1 reflects the models’ fundamental conceptual reasoning abilities, specifically their capacity to correctly determine the truth or falsehood of a given statement. While open-source models exhibit some performance, their overall performance is relatively low around 30. Even the advanced Qwen-2.5-Math-72B-Instruct achieves only 41.8, falling behind commercial models. Notably, the math-reasoning-specific o1 model outperforms others, achieving an F1 score of 60.1. However, compared to benchmarks focused on elementary and high school mathematics, the overall performance is significantly lower. This disparity underscores the challenges posed by our higher-level mathematics benchmark, aligning with our objective of exploring deeper conceptual reasoning. When comparing the performance of models of different sizes, we find that some small 7B models outperform the larger 72B model. For example, Deepseek-Math-7B-rl outperforms WizardMath-70B-v1.0. We believe this phenomenon is related to the training methods behind these models. As explained in Section 1, drill-based training in WizardMath-70B-v1.0 is not very effective in improving the ability to understand deep and complex mathematical concepts. In contrast, the superior performance of Deepseek-Math-7B-rl suggests that reinforcement learning may play a key role in enhancing counterexample-driven conceptual reasoning of LLMs.
@@
Models GSM8K MATH
GPT-4o-2024-08-06 92.9 81.1
Qwen2.5-math-7B-Instruct 95.1 80.5
Qwen2.5-math-72B-Instruct 95.4 84.9
@@ 6.2 Results with Finetuning
To further validate the effectiveness of counterexample reasoning for conceptual reasoning, we construct a training dataset designed for counterexample-based reasoning, conduct training, and evaluate the performance, as shown in Table 2 and Table 3.
@@
We evaluate the approach on our benchmark to verify whether counterexample reasoning effectively enhances a model’s conceptual reasoning capabilities. In addition to the base model, we include comparisons with the hint prompt, where explicit hints encouraging LLM to reason by example were provided. For a fair comparison, we use identical prompts to generate outputs during the experiments. The results indicate that, with just 1,025 training samples, the fine-tuned model outperforms all base models in judgement F1 score. Furthermore, the trained model demonstrates superior example-based conceptual reasoning abilities, showing improvements in both the quantity and quality of examples compared to base models. On the other hand, although the constructed training data are refined to align closely with our CounterMATH distribution, some discrepancies remain. Since our exploration involves only the simple SFT strategy on a limited dataset, the model performs slightly worse on some metrics, which can be considered an acceptable limitation.
@@
To assess the generalizability of the fine-tuned model, we further evaluate its performance on out-of-distribution (OOD) benchmarks. This aims to verify whether the model’s counterexample reasoning capability, which is valid on our CounterMATH, can transfer to other benchmarks and deliver broader performance improvements. Using identical prompts and configurations for fairness, we compare the base models and fine-tuned models on OOD benchmarks MATH and GSM8K. The results reveal that the fine-tuned model outperforms the base models on both benchmarks, even surpassing larger models like the 72B-parameter model. This aligns well with our hypothesis: equipping the model with counterexample reasoning ability can enhance conceptual reasoning across the general mathematical domain.
@@ Appendix B Details for Experimental Settings
Please reason step by step about whether the above statement is True or False, and put your final answer within $\backslash\backslash\ \boxed{}$.
@@
Please reason step by step about whether the above statement is True or False, and put your final answer within $\backslash\backslash\ \boxed{}$.
@@
Please reason step by step about whether the statement is True or False, and put your final answer within $\backslash\backslash\ \boxed{}$.<EOS>
@@
Please reason by giving examples about whether the above statement is True or False, and put your final answer within $\backslash\backslash\ \boxed$.

### v2

@@ 1 Introduction
As illustrated in Figure 1, for human mathematics learning, “example-based” learning is a more important strategy than drill-based learning. In particular, for mathematical proofs, “proof by counterexamples” is an indispensable approach. Inspired by the idea that counterexample-driven proofs better reflect a deep understanding of mathematical concepts, we propose CounterMATH  [^1] https://github.com/THUKElab/COUNTERMATH, a counterexample-based mathematical reasoning benchmark. CounterMATH is designed to evaluate LLMs’ ability to distinguish subtle differences between mathematical terms and properties at university-level by providing examples. Specifically, we collect 1,216 statement-rationale pairs from mathematical textbooks, focusing on disproving certain statements under unusual conditions using counterexamples. In terms of difficulty, CounterMATH covers advanced mathematical knowledge similar to PutnamBench (Tsoukalas et al., ) and Putnam-AXIOM (Gulati et al., 2024), both of which assess the depth of mathematical understanding in LLMs.
@@ 4 Benchmark Settings
We are testing following large language models with CounterMATH. For open-weight models, we use Deepseek-Math-7B-RL (Shao et al., 2024), Eurus-2-7B-PRIME (Cui et al., 2025), Qwen2.5-Math-7B/72B-Instruct (Yang et al., 2024), NuminaMath-7B-TIR (Li et al., 2024b), InternLM2-Plus-7B/20B/Mixtral8x22B (Ying et al., 2024), Abel-7B/13B/70B (Chern et al., 2023), WizardMath-7B/70B (Luo et al., 2023), Mathstral-7B [^4] https://mistral.ai/news/mathstral/, MetaMath-Mistral-7B (Yu et al., 2023), Xwin-Math-7B/13B/70B (Li et al., 2024a), Rho-math-7b-interpreter (Lin et al., 2024b), MAmmoTH2-7B/8x7B-Plus (Yue et al., 2024), QwQ-32B-Preview [^5] https://qwenlm.github.io/blog/qwq-32b-preview/. For proprietary models, we use GPT-4o, OpenAI o1666https://cdn.openai.com/o1-system-card-20241205.pdf, Qwen-Max [^7] https://www.alibabacloud.com/help/en/model-studio/developer-reference/what-is-qwen-llm, Deepseek-R1888https://api-docs.deepseek.com/news/news250120, Claude3.7-sonnet [^9] https://www.anthropic.com/claude/sonnet, and Gemini2.5-pro [^10] https://deepmind.google/technologies/gemini/pro/. The selection of baseline models tries to cover as many representative models as possible, which considers various perspectives, including data processing, training paradigms, base models, and developers (i.e., academia or industry). The detailed summary of open-weight baseline models is in Appendix B.
@@
5 Conceptual Finetuning
5.1 Training Data Engineering Framework
Filter-based Data Collection
@@ 5.1 Training Data Engineering Framework
Deepseek-R1 80.7 86.8 54.2 65.3
Claude3.7-sonnet 64.8 78.0 45.0 52.5
Gemini2.5-pro 77.0 90.8 65.1 75.7
@@ 5.2 Training and Evaluation Details
For model training, we select Qwen-2.5-Math-7B-Instruct, an open-source model known for its strong mathematical reasoning capabilities and general applicability. we fine-tune Qwen2.5-Math-Instruct-7B using supervised LoRA training on $\timesL2$0 48GB GPUs, with a learning rate of 1.0e-5. After training, we evaluate the model on both our CounterMATH and several out-of-distribution (OOD) benchmarks, such as MATH(Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021), OlympaidBench (He et al., 2024), MMLU (Wang et al., 2024c), AIME 2024. We aim to demonstrate not only improved performance on the CounterMATH but also enhanced generalization to OOD benchmarks, thereby validating the effectiveness of our insight in improving the model’s overall conceptual reasoning capabilities by examples.
@@ 6.1 Evaluation Results without Finetuning
The performance on the automatic evaluation metric F1 reflects the models’ fundamental conceptual reasoning abilities, specifically their capacity to correctly determine the truth or falsehood of a given statement. While open-source models exhibit some performance, their overall performance is relatively low around 30. Even the advanced Qwen-2.5-Math-72B-Instruct achieves only 41.8, falling behind commercial models. Notably, the math-reasoning-specific Deepseek-R1 model outperforms others, achieving an F1 score of 80.7. However, compared to benchmarks focused on elementary and high school mathematics, the overall performance of commercial models is relatively lower. This disparity underscores the challenges posed by our higher-level mathematics benchmark, aligning with our objective of exploring deeper conceptual reasoning. When comparing the performance of models of different sizes, we find that some small 7B models outperform the larger 72B model. For example, Deepseek-Math-7B-rl outperforms WizardMath-70B-v1.0. We believe this phenomenon is related to the training methods behind these models. As explained in Section 1, drill-based training in WizardMath-70B-v1.0 is not very effective in improving the ability to understand deep and complex mathematical concepts. In contrast, the superior performance of Deepseek-Math-7B-rl suggests that reinforcement learning may play a key role in enhancing counterexample-driven conceptual reasoning.
@@
Models GSM8K MATH Olympaidbench-Math MMLU-college-Math AIME 2024
Qwen2.5-math-7B-Instruct 95.1 80.5 41.6 74.0 20.0
Qwen2.5-math-72B-Instruct 95.4 84.9 49.2 78.3 26.7
@@ 6.2 Results with Finetuning
We evaluate the approach on our benchmark to verify whether counterexample reasoning effectively enhances a model’s conceptual reasoning capabilities. In addition to the base model, we include comparisons with the hint prompt, where explicit hints encouraging LLM to reason by example were provided. For a fair comparison, we use identical prompts to generate outputs during the experiments. The results indicate that, with just 1,025 training samples, the fine-tuned model outperforms base models in F1 score. Furthermore, the trained model demonstrates superior example-based conceptual reasoning abilities, showing improvements in both the quantity and quality of examples compared to base models. On the other hand, although the constructed training data are refined to align closely with our CounterMATH distribution, some discrepancies remain. Since our exploration involves only the simple SFT strategy on a limited dataset, the model performs slightly worse on some metrics, which can be considered an acceptable limitation.
@@
To assess the generalizability of the fine-tuned model, we further evaluate its performance on out-of-distribution (OOD) benchmarks. This aims to verify whether the model’s counterexample reasoning capability, which is valid on our CounterMATH, can transfer to other benchmarks and deliver broader performance improvements. Using identical prompts and configurations for fairness, we compare the base models and fine-tuned models on OOD benchmarks. The results reveal that the fine-tuned model outperforms the base models on both benchmarks, even surpassing larger models like the 72B-parameter model. This aligns well with our hypothesis: equipping the model with counterexample reasoning ability can enhance conceptual reasoning across the general mathematical domain.
@@ 6.3 Used Tokens Analysis
Figure 7: Our case studies of the performance of LLM include: (1) correct judgment with incorrect examples, (2) validation of novel but effective counterexamples, and (3) performance improvement on MATH dataset after counterexample training
6.4 Case Study
In our case study, we observe that while the LLM correctly judges mathematical statements, it sometimes provides incorrect examples as shown in Figure 7, Case 1. In addition, we verify the validity of our evaluation metric of EXAMPLE. The LLM generates valid counterexamples that differs from reference ones, yet are correctly validated by our judge model, highlighting the effectiveness of our evaluation method. Finally, we compare the improvement of MATH benchmark between base and SFT model, demonstrating the benefit of incorporating counterexample traning to enhance mathematical reasoning capabilities.
@@ 7 Conclusion
Acknowledgements
This research is supported by National Natural Science Foundation of China (Grant No. 62276154), Research Center for Computer Network (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No. 2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No. JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006), the Major Key Project of PCL (NO. PCL2024A08), the Key-Area Research and Development Program of Guangdong Province No.2024B1111100001. This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
@@ Appendix B Details for Experimental Settings
Please reason step by step about whether the above statement is True or False, and put your final answer within $\backslash\backslashboxed{}$.
@@
Please reason step by step about whether the above statement is True or False, and put your final answer within $\backslash\backslashboxed{}$.
@@
Please reason step by step about whether the statement is True or False, and put your final answer within $\backslash\backslashboxed{}$.<EOS>
@@
Please reason by giving examples about whether the above statement is True or False, and put your final answer within $\backslash\backslashboxed$.